OpenStudy (anonymous):

Find a number whose product with 4 is the same as its sum with 12 ?

OpenStudy (anonymous):

let that number be "x" so according to the given condition 4x = 12 + x can you solve it now..??

OpenStudy (anonymous):

@StudentAthlete ???

OpenStudy (anonymous):

so it is 3 ?

OpenStudy (anonymous):

no... 4x = 12 + x subtract "x" on both sides 4x-x = 12 + x - x 3x = 12 divide by "3" on both sides 3x/3 = 12/3 x = 4

OpenStudy (anonymous):

@StudentAthlete got it Now..??

OpenStudy (anonymous):

yess thanks